HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

TUNING THE METHYL TORSION IN 2-METHYLTHIOPHENE DERIVATIVES BY RING SUBSTITUTION

\underline{K. J. KOZIOL} ^{1}, H. V. L. NGUYEN^{1,2},^{1}Univ Paris Est Créteil and Université Paris Cité, CNRS, LISA, 94010 Créteil, France,^{2}Institut Universitaire de France

(IUF), 1 Rue Descartes, 75231 Paris, France

2-methylthiophene (2MTP) features a methyl group at the 2-position of the thiophene ring that undergoes internal rotation hindered by a V_{3} barrier of 197.6 cm^{−1}.^{1} Adding another substituent at the 5-position has a major impact on the barrier height, which appeared to be of electronic origin due to the absence of sterical hindrance. If a chlorine atom is present at the 5-position, as in 2-chloro-5-methylthiophene (2Cl5MTP), the V_{3} barrier increases to 246.2 cm^{−1}. Being separated by the sulfur atom, the apparent influence of chlorine on the methyl internal rotation can only be explained by electronic properties transferred through the π-conjugated system. While it is surprising that replacing the chlorine atom that has a strong −I and +M effect with a second methyl group, as in 2,5-dimethylthiophene, leads to a minor change in the V_{3} barrier (248.0 cm^{−1}),^{2} replacing it with an acetyl group as in syn-2-acetyl-5-methylthiophene greatly reduced the barrier to 157.3 cm^{−1}.^{3} Comparison of the barriers to those of the corresponding furan derivates shows a large impact of the hetero atom with an increase to 412.9 cm^{−1} for 2-methylfuran,^{4} 439.1 cm^{−1} for 2,5-dimethylfuran,^{5} and 356.5 cm^{−1} for syn-2-acetyl-5-methylfuran.^{6} The methyl torsional barrier of 2-methylfuran derivates is tuned by the same manner as it is for the thiophene analogues. For 2MTP and 2Cl5MTP, we used a state-of-the-art pulsed molecular jet Fourier transform spectrometer to measure the microwave spectra in the frequency range of 2-26.5 GHz. The spectra of the ^{34}S and ^{13}C isotopologues of 2MTP could be observed in natural abundances and the semi-experimental r_{e}^{SE} structure was determined. for 2Cl5MTP, the ^{35}Cl and ^{37}Cl nuclei give rise to a nuclear quadrupole hyperfine structure in the spectrum in addition to the A-E torsional splittings of the methyl internal rotation.

1doi:10.3390/spectroscj1010005, K. J. Koziol, H. El Hadki, A. Lüchow, N. Vogt, J. Demaison, H. V. L. Nguyen, Spectrosc. J. 1, 49-64 (2023).

2doi:10.1039/C5CP03513A, V. Van, W. Stahl, H. V. L. Nguyen, Phys. Chem. Chem. Phys. 17 32111-32114 (2015).

3doi:10.1039/D2CP03897H, C. Dindic, H. V. L. Nguyen, Phys. Chem. Chem. Phys. 25, 509-519 (2023).

4doi:10.1016/j.jms.2012.06.005, I. A. Finneran, S. T. Shipman, S. L. W. Weaver, J. Mol. Spectrosc.

280, 27-33 (2012).

5doi:10.1016/j.jms.2017.11.007, V. Van, J. Bruikhuisen, W. Stahl, V. Illyushin, H. V. L. Nguyen, J.

Mol. Spectrosc. 343, 121-125 (2018).

6doi:10.1002/cphc.201600757, V. Van, W. Stahl, H. V. L. Nguyen, Chem. Phys. Chem. 17, 3223-3228 (2016).

p-number: p162  Submitted on Wed, 14 Jun 2023 23:59:04 +0200